Given that Kw for water is 2.4×10−14 at 37C, calculate the pH of a neutral aqueous solution at 37C, which is the normal human body temperature
1
Expert's answer
2020-04-18T06:45:35-0400

Kw=[H+]*[OH-];

If the aqueous solution is neutral, [H+]=[OH-], then

[H+]=[OH-]=√Kw=√2.4*10^-14 = 1.55×10^-7

pH=-lg[H+]=6.8